The solutions to the inequality ( [(x1/2)(x-5)] / lx-7l )>0 are all the x elements of R such that?


A. x>0 and x


1
Expert's answer
2021-08-11T19:22:19-0400
"\dfrac{x^{1/2}(x-5)}{|x-7|}>0"

"x^{1/2}: x\geq0"

"|x-7|\not=0=>x\not=7"

Then


"x-5>0, x\not=0, x\not=7"

The solution is


"x\in(5, 7)\cup(7, \infin)"

The statement that the solutions to the inequality "\dfrac{x^{1/2}(x-5)}{|x-7|}>0" are all the "x" elements of "\R"  is False.